Find the ∆E



a) When 175J of work is done on a system that evolves 50J of heat to the surrounding. b) When a gas absorbs 28kJ of heat and 13kJ of work done on it.

1
Expert's answer
2022-03-29T05:09:03-0400

∆U =change in internal energy= Q + W

∆U=?

(a) Given conditions

W= +175J

Because work is done on the system

Q= - 50J

Because heat release from the system to the surrounding

∆U =change in internal energy= Q + W

∆U=?

∆U= -50 J + 175 J = 125 J Ans.

(b) Q = + 28 KJ

Because gas absorbs heat to the system from the surrounding

W = + 13 KJ

Because work is done on the system

∆U =change in internal energy= Q + W

∆U = 28 KJ + 13 KJ

∆U = 41 KJ = 41000 J And.